OpenStudy (anonymous):

help please? The sum of two numbers is 36. Twice the first number minus the second is 6. Find the numbers. A. (10, 26) B. (30, 6) C. (14, 22) D. (42, 6)

OpenStudy (anonymous):

The first number is 14 and the second is 22.

OpenStudy (anonymous):

C

OpenStudy (anonymous):

a+b=36 2a-b=6 -------- cancel b a+2a=36+6 3a=42 a=14 b=22

OpenStudy (primeralph):

@mz_little10 @nickersia really, stop posting answers. You're not helping

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

No problem :)